A company is projected to generate a free cash flow of $100 million next year (year 1) and $120 million in two years (year 2). After that it is projected grow at a steady rate in perpetuity. The company's cost of capital is 10%. It has $400 million of debt and $40 million in cash. There are 60 million shares outstanding. Comparable companies trade at an average EV/FCFF multiple of 6.9. Using the exit multiple method for terminal value and DCF for the rest, what is your estimate of its share price? Round to one decimal place ​[Hint: Draw the timeline. Compute TV2 = Multiple x FCFF2. Discount FCFF1, FCFF2 and TV2. Walk the bridge to stock price]
Answer #1

Solution) Free cash flow to firm for year 1 (FCFF1) = $100 million

Free cash flow to firm for year 2 (FCFF2) = $120 million

Third-year onwards, the FCFF grows at a constant rate till perpetuity.

EV/FCFF multiple = 6.9

The value of this perpetuity can be estimated at the end of year 2 using the EV/FCFF multiple

Terminal Value at the end of year 2 (TV2) = EV/FCFF multiple*FCFF2 = 6.9*120 = $828 million

Cost of capital = 10%

The Cash Flows are as follows:

t=0 t=1 t=2
FCFF $        100 $        120
TV2 $        828
Total Cash Flows $        100 $        948
Enterprise Value =NPV(cost of capital, cash flows)
Enterprise Value $ 874.38 million

Value of Debt = $400 million

Cash & Cash Equivalents = $40 million

Enterprise Value (EV) = Market Value of Equity + Market Value of Debt - Cash & Cash Equivalents

874.38 = Market Value of Equity + 400 - 40

Market Value of Equity = 874.38 - 400 + 40 = $514.38 million

Market Value of Equity = Price per share * Number of outstanding shares

Price per share = Market Value of Equity/Number of outstanding shares

Number of outstanding shares = 60 million

Price per share = 514.38/60 = 8.573